The diagonal, d of a box can be found using the formula d = where l, w, and h , represent the length, width and height of the box. If the box is 24 cm in length, 8 cm in width, and 6 cm in height, then what is the length of the diagonal?
Answer by Edwin McCravy(20077)




You inadvertently left out the formula, so I'll do it without the formula
first and then again with a formula.  

We want the length of the red line DB, which is the
diagonal of the box.  

There are two right triangles ABC and DBC.

First we calculate the hypotenuse BC of right triangle
ABC using the Pythagorean theorem:

 BC² = AC² + AB²
 BC² = 8² + 24²
 BC² = 64 + 576
 BC² = 640
 BC = √640
 BC = 8√10

The required diagonal, the red line DB is the hypotenuse of
right triangle DBC.  So we use the Pythagorean theorem again

DB² = DC² + BC²
DB² = 6² + 640   (Notice we only needed BC², not BC)
DB² = 36 + 640
DB² = 676
 DB = √676
 DB = 26

So the diagonal DB is 26 cm in length, without using the formula,
but only the Pythagorean theorem.

-----------------------------------------------

The formula you omitted is:

d = √L² + W² + H² 

Substituting gives

d = √L² + W² + H²

d = √24² + 8² + 6²

d = √576 + 64 + 36

d = √676

d = 26 cm

Edwin